OpenStudy (anonymous):

Which number equals the square root of 1, 029.? 7the square root of 21. 14the square root of 3. 3the square root of 21. 21the square root of 3.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: 1029 = 49*21

OpenStudy (anonymous):

Ooh, thanks.. I'll try to go from there and I'll tag you if I get stuck.

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

I feel like it's the first one or the middle one, or am I wrong?

jimthompson5910 (jim_thompson5910):

did you factor and break up the square root

OpenStudy (anonymous):

Oh, I had a feeling I should factor more. So the ending would be a 3, for sure, right?

jimthompson5910 (jim_thompson5910):

no \[\large \sqrt{1029} = \sqrt{49*21}\] \[\large \sqrt{1029} = \sqrt{49}*\sqrt{21}\] keep going

OpenStudy (anonymous):

Oh, jeez I'm dumb, okay, one sec.

jimthompson5910 (jim_thompson5910):

you're not dumb

OpenStudy (anonymous):

You're right, I'm not. I'm only dumb with Algebra 1. Is the answer the first one?

jimthompson5910 (jim_thompson5910):

you just need practice, that's all

jimthompson5910 (jim_thompson5910):

yes \[\large \sqrt{1029} = 7\sqrt{21}\]

OpenStudy (anonymous):

YAY I actually got it for once. :) And yeah, these are my last few lessons for this year for ALG1 and then on to Geometry.

jimthompson5910 (jim_thompson5910):

nice work

jimthompson5910 (jim_thompson5910):

and glad you got it